OpenStudy (anonymous):

Find the Coordinate and the slope. y-5=3 x (Show Work)

OpenStudy (anonymous):

-5 from both sides y=3x-5 y=mx+b m=3/1 or 3 b=y-intercept

OpenStudy (anonymous):

but what happnes to the x

OpenStudy (anonymous):

the x just helps you set up the eqation in that form and I forgot m=slope

OpenStudy (anonymous):

ok but wouldnt you add 5 to both sides cuz its oppisite

OpenStudy (anonymous):

sorry yes I made a mistake you are correct

OpenStudy (anonymous):

ok and thats it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and whats the coordinate

OpenStudy (anonymous):

I'm sorry the coordinate is any point that passes through the equation when graphed. http://www.mathway.com/answer.aspx?p=grap?p=ySMB013x+5?p=True?p=True?p=True?p=True?p=True

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

no prob